# Overcoming Reward Overoptimization via Adversarial Policy Optimization with Lightweight Uncertainty Estimation

### v1

@@ Appendix C Theoretical Results
Let $\|\phi^{*}-\hat{\phi}\|_{M}\leq\sqrt{B}$. For adversarial search with the max-min objective, it relaxes $A_{1}$ through:
$\left|\mathbb{E}_{x,y\sim\pi_{\theta}(\cdot|x)}\left[g_{x,y}^{T}(\phi^{*}-\hat{\phi})\right]\right|\leq\left|\mathbb{E}_{x,y\sim\pi_{\theta}(\cdot|x)}\left[g_{x,y}\right]^{T}(\phi^{*}-\hat{\phi})\right|\leq\sqrt{B}\cdot\|\mathbb{E}_{x,y\sim\pi_{\theta}(\cdot|x)}\left[g_{x,y}\right]\|_{M^{-1}}.$
@@
$\left|\mathbb{E}_{x,y\sim\pi_{\theta}(\cdot|x)}\left[g_{x,y}^{T}(\phi^{*}-\hat{\phi})\right]\right|\leq\mathbb{E}_{x,y\sim\pi_{\theta}(\cdot|x)}\left[|g_{x,y}^{T}(\phi^{*}-\hat{\phi})|\right]\leq\sqrt{B}\cdot\mathbb{E}_{x,y\sim\pi_{\theta}(\cdot|x)}\left[\||g_{x,y}\|_{M^{-1}}\right].$ (15)
@@ Appendix D Details on Gaussian Processes and Bayesian linear regression
D.1 Bayesian Linear regression for Uncertainty Quantification
In BLR, the projection weights $\phi$ are presumed to follow a Gaussian prior, expressed as $p(\phi)=\mathcal{N}(\phi|\mathbf{m}_{0},\mathbf{S}_{0})$, with $\mathbf{m}_{0}$ being the prior mean and $\mathbf{S}_{0}$ the prior covariance matrix.
Then the predictive posterior mean $\mu_{\text{post}}$ and variance $\sigma^{2}_{\text{post}}$ of the reward for a new datapoint $(x,y)$ in BLR can be expressed in closed form as follows:
$p(r|x,y,\mathcal{D})$ $=\mathcal{N}(r|\mu_{\text{post}},\sigma^{2}_{\text{post}})$ (16)
$\mu_{\text{post}}$ $=e(x,y)^{T}\mathbf{m}_{D},$ (17)
$\sigma_{\text{post}}$ $=\sqrt{e(x,y)^{T}\mathbf{S}_{D}^{-1}e(x,y)+\epsilon^{2}}.$ (18)
where:
$\mathbf{S}_{N}^{-1}$ $=\mathbf{S}_{0}^{-1}+\Phi^{T}\Phi/\epsilon^{2},$
$\mathbf{m}_{N}$ $=\mathbf{S}_{N}(\mathbf{S}_{0}^{-1}\mathbf{m}_{0}+\Phi^{T}\mathbf{r}/\epsilon^{2}),$
where $\mathbf{S}_{D}$ and $\mathbf{m}_{D}$ are the closed form posterior mean and variance of the projection weights $\phi$ calculated based on the reward training data $D$. $\Phi$ is the design matrix formed by stacking the feature vectors $e(x_{i},y_{i})$, $\mathbf{r}$ is the vector of outputs $r_{\varphi}(x_{i},y_{i})$ and $\epsilon$ being the aleatoric noise.
D.2 Gaussian Processes for Uncertainty Quantification
@@
where $\Phi$ denotes the matrix of training input embeddings i.e. the same as in BLR, $\mathbf{r}$ is the vector of training outputs, $K$ represents the covariance matrix computed using the kernel function $k$ over the training embeddings, $\sigma^{2}_{n}$ signifies the noise variance, and $I$ is the identity matrix. In practice, the Radial Basis Function (RBF) kernel, also known as the Gaussian kernel, is the most popular choice due to its flexibility and the property of being a universal kernel. The RBF kernel is defined as:
@@
we recover Bayesian Linear Regression (BLR). This is because the linear kernel implies a linear relationship between the inputs, consistent with the assumptions of BLR. In all our experiments, we used 2000 randomly sampled training datapoints to construct the GP uncertainties.

### v2

@@ Appendix C Theoretical Results
Recall that $\|\phi^{*}-\hat{\phi}\|_{M_{D}}\leq b$. For adversarial search with the max-min objective, it relaxes $A_{1}$ through:
$\left|\mathbb{E}_{x,y\sim\pi_{\theta}(\cdot|x)}\left[g_{x,y}^{T}(\phi^{*}-\hat{\phi})\right]\right|\leq\left|\mathbb{E}_{x,y\sim\pi_{\theta}(\cdot|x)}\left[g_{x,y}\right]^{T}(\phi^{*}-\hat{\phi})\right|\leq b\cdot\|\mathbb{E}_{x,y\sim\pi_{\theta}(\cdot|x)}\left[g_{x,y}\right]\|_{M^{-1}}.$
@@
$\left|\mathbb{E}_{x,y\sim\pi_{\theta}(\cdot|x)}\left[g_{x,y}^{T}(\phi^{*}-\hat{\phi})\right]\right|\leq\mathbb{E}_{x,y\sim\pi_{\theta}(\cdot|x)}\left[|g_{x,y}^{T}(\phi^{*}-\hat{\phi})|\right]\leq b\cdot\mathbb{E}_{x,y\sim\pi_{\theta}(\cdot|x)}\left[\||g_{x,y}\|_{M^{-1}}\right].$ (16)
@@ Appendix D Details on Gaussian Processes and Bayesian linear regression
D.1 Gaussian Processes for Uncertainty Quantification
@@
where $\Phi$ denotes the matrix of training input embeddings i.e. $(e(x_{1},y_{1}),...,e(x_{n},y_{n}))$, $k(e(x^{*},y^{*}),\Phi):=(k(e(x^{*},y^{*}),e(x_{1},y_{1})),\dots k(e(x^{*},y^{*}),e(x_{n},y_{n})))^{T}$, $\mathbf{r}$ is the vector of training outputs, $K$ represents the covariance matrix computed using the kernel function $k$ over the training embeddings, $\sigma^{2}_{n}$ signifies the noise variance, and $I$ is the identity matrix. In practice, the Radial Basis Function (RBF) kernel, also known as the Gaussian kernel, is the most popular choice due to its flexibility and the property of being a universal kernel. The RBF kernel is defined as:
@@
we recover Bayesian Linear Regression (BLR). This is because the linear kernel implies a linear relationship between the inputs, consistent with the assumptions of BLR. In all our experiments, we used 2000 randomly sampled training datapoints to construct the GP uncertainties. Below, we add experiments demonstrating that GP have similar ability to detect overoptimization.
Figure 3: Anthorpic HH Dataset: Comparison among lightweight uncertainty estimations. (left) the blue lines with shaded areas depict the reward dynamics concerning optimization steps in PPO, where the solid and dashed lines represent gold and proxy rewards, respectively. The lines with uncertainty bars are results from different uncertainty estimation methods. The reward values are indexed on the left y-axis, while the uncertainty is indexed on the right y-axis. (Right) We plot the correlation between uncertainty and the difference between gold rewards and proxy rewards.
Figure 4: TL;DR Dataset: Comparison among lightweight uncertainty estimations. (left) the blue lines with shaded areas depict the reward dynamics concerning optimization steps in PPO, where the solid and dashed lines represent gold and proxy rewards, respectively. The lines with uncertainty bars are results from different uncertainty estimation methods. The reward values are indexed on the left y-axis, while the uncertainty is indexed on the right y-axis. (Right) We plot the correlation between uncertainty and the difference between gold rewards and proxy rewards.
The above figures show that GP uncertainty estimates also correlate with the increase of the difference between the estimated and golden reward (right Figures). On the Anthropic HH dataset GP seems to be similar to our CI, however on the TL;DR Dataset we note on the left Figure 4 that the uncertainty of GP increases significantly from step $100-500$, whereas it does not for CI. In fact for step $100-500$, the uncertainty should be small as suggested by our CI, because the predicted reward and golden reward are indeed similar. Therefore we mainly opted for the CI method in our paper and leave this interesting direction of using Bayesian methods to future work.